OpenStudy (anonymous):

Suppose that y varies directly with x and inversely with z, and y = 18 when x = 15 and z = 5. Write the equation that models the relationship. Then find y when x = 21 and z = 7

OpenStudy (anonymous):

\[y=C \times \frac{x}{z}\]Now use your initial conditions to find C, and then use that C to find the second y point.

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

Since y varies directly with x, as x gets bigger so does y. As x gets smaller, so does y. That is why in the equation I wrote, I multiplied C by x. The same reason has me divide C by z. Since as z gets smaller y gets bigger and as z gets bigger y gets smaller, we know that we must be dividing something by z. So since when x=15 and z =5, y=18, I can say\[y=C \times \frac{x}{z}\]\[18=C \times \frac{15}{5}= C \times 3\]\[C=6\]So I can say\[y=6 \times \frac{x}{z}\]Now all you have to do is plug in the new values for x and z.

OpenStudy (anonymous):

would it be t = 6x/z; 18 or t = 6z/x; 2

OpenStudy (anonymous):

I've basically already given you the answer.

OpenStudy (anonymous):

oh. thanks!!